A die (20 sided with numbers 1-20) is rolled 650 times. Find the standard deviation for the number of "11"s that will be rolled.
Answer by stanbon(75887)

A die (20 sided with numbers 1-20) is rolled 650 times. Find the standard deviation for the number of "11"s that will be rolled.
---------
P(x = 11) = 1/20
----
n = 650
----
q = 19/20
------
std = sqrt[npq] = sqrt[650(1/20)(19/20)] = (1/20)sqrt(650*19) = 5.557..
Cheers,
Stan H.
============